For the function f(x)= x^2-2 defined over[1,6],





Verify L(P,f)≤U(P,f) where is the partition which divide [1,6] into five equal intervals

1
Expert's answer
2022-01-14T06:48:30-0500

If [1,6] is divided into 5 equal partitions, the partitions will be :-

"P=\{P_i\}=\{[x_i,x_{i+1}]\}=\{[i,i+1]\}, \\(i=1,2,3,4,5)\\
\text{Now we know that,}\\
M_i= sup\{f(x) : x_i\leq x\leq x_{i+1} \}\\
=sup\{f(x) : i\leq x\leq {i+1} \}\\
\text{And, }\\
m_i= inf \{f(x) : x_i\leq x\leq x_{i+1} \}\\
=inf\{f(x) : i\leq x\leq {i+1} \}"

As f(x) is a increasing function over positive numbers so,

"M_i=f(x_{i+1})=f(i+1)=(i+1)^2-2\\
=i^2+2i-1\\
\text{and, }\\
m_i=f(x_i)=f(i)=i^2-2
\\\text{Now, }\Delta x_i=x_{i+1}-x_i=i+1-i=1"

"\therefore U(P,f)-L(,U,f)\\
=\sum_{i=1}^5 M_i\Delta x_i-\sum_{i=1}^5 m_i\Delta x_i\\
=\sum_{i=1}^5 (M_i-m_i)\Delta x_i\\
=\sum_{i=1}^5 (i^2+2i-1-i^2+2)\times 1\\
=\sum_{i=1}^5 (2i+1) \\= 2\times \dfrac{5\times 6}{2}+5=35>0
\\\Rightarrow U(P,f)>L(P,f)"

(Hence verified.)